Note:

1 mole of a compound is equivalent to its molecular mass

1 mole of a compound is equal to 6.02x10^23 molecules

The percentage is a direct equivalent to grams

Answers should be in 3 significant figures

No need to include the units and the compound in the answers


  1. A compound is found to contain 63.52 % iron and 36.48% sulfur. Find its empirical formula. 
  2.   A compound is found to contain 26.56% potassium, 35.41% chromium, and the remainder oxygen. Find its empirical formula.  

 

  




1
Expert's answer
2021-06-29T00:38:37-0400

1. One gram of a compound contain 0.6352 g iron and 0.3648 g sulfur.

"n(Fe) = \frac{0.6352}{55.84} = 0.011375 \; mol \\

n(S) = \frac{0.3648}{32.06} = 0.011378 \; mol"

Proportion:

n(Fe) : n(S) = 0.011375 : 0.011378 = 1:1

Empirical formula: FeS

2. One gram of a compound contain 0.2656 g potassium and 0.3541 g chromium, and 0.3803 g oxygen.

"n(K) = \frac{0.2656}{39.10}=0.006792 \;mol \\

n(Cr) = \frac{0.3541}{51.99}= 0.006810 \;mol \\

n(O) = \frac{0.3803}{15.99}= 0.023783 \;mol"

Proportion:

n(K) : n(Cr) : n(O) = 0.006792 : 0.006810 : 0.023783

"= \frac{0.006792}{0.006792} : \frac{0.006810}{0.006792} : \frac{0.023783}{0.006792} \\

= 1 : 1 : 3.5 = 2: 2: 7"

Empirical formula: K2Cr2O7